# FaaF: Facts as a Function for the evaluation of RAG systems

### 06 Mar 2024 v1

@@ Abstract
Factual recall from a reference source is crucial for evaluating the performance of Retrieval Augmented Generation (RAG) systems, as it directly probes into the quality of both retrieval and generation. However, it still remains a challenge to perform this evaluation reliably and efficiently. Recent work has focused on fact verification via prompting language model (LM) evaluators, however we demonstrate that these methods are unreliable in the presence of incomplete or inaccurate information. We introduce Facts as a Function (FaaF), a new approach to fact verification that utilizes the function calling abilities of LMs and a framework for RAG factual recall evaluation. FaaF substantially improves the ability of LMs to identify unsupported facts in text with incomplete information whilst improving efficiency and lowering cost by several times, compared to prompt-based approaches.
@@ 1 Introduction
The adoption and transformative impact of large language models (LMs) across industries are significantly propelled by their application in knowledge-intensive tasks.
@@
Figure 1: An overview of FaaF, a constructor dynamically creates a function object from a set of fact statements. Function calling allows LMeval to verify all facts within a single call when provided with an input text. FaaF reduces the error rate in identifying unsupported facts by up to 40 percentage points compared to prompting whilst reducing the number of LMeval calls and output tokens by more than 5 times.
@@
Retrieval Augmented Generation (RAG) enables seamlessly the integration of out-of-training knowledge to the LM (Lee et al., 2019). This is achieved by dynamically fetching and incorporating relevant data from a trusted reference source prior LM text generation, resulting in more informed, accurate, and contextually rich outputs. In such applications, factual recall —the extent to which all the information in the reference source that is required to sufficiently answer the question is retrieved and correctly conveyed in the response of the RAG— is a key performance metric because it closely reflects the central purpose of these systems and captures the the performance of retrieval and generation simultaneously.
Although there is a considerable body of work (Chen et al., 2022; Zhang et al., 2023; Gao et al., 2023; Lee et al., 2023; Min et al., 2023; Azaria and Mitchell, 2023; Yuan et al., 2021; Fu et al., 2023; Wang et al., 2023; Kadavath et al., 2022; Manakul et al., 2023) that is focusing at factual precision —the truthfulness of each statement in a generated text— factual recall at least as important for RAG systems, where a response can be factually precise but given the wrong context may be irrelevant to the posed question.
Practical factual recall evaluation in RAG systems is still an open problem because of the following reasons. First, the scenario of automatic verification of truthful, independent statements, given a highly incomplete or inaccurate generated text (simulating recall in a low quality RAG) has not been as explored as precision on LM generated text. Secondly, considering that there can be several facts which need to be verified in a single generated text (E.g. a RAG response to a question), current approaches relying on verification of each fact independently can be prohibitively costly in time and resources. RAG systems include many moving parts (knowledge base, retrieval, prompt formulation, LM) and require substantial tuning (Es et al., 2023) therefore efficiency is a requirement for practical use.
@@
Other recent works (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) attempt to evaluate LMs and RAG systems in terms of factual recall by exact matching of ground truth text in the generated text. While this is a fast and efficient process, it faces serious challenges and limitations in the verification task. As recognised by Cuconasu et al. (2024), exact matching of ground truth text in the generated text is prone to false negatives since the ground truth information might be present in the generated text but phrased differently. Secondly, this approach cannot be applied when the ground truth information longer than a few words since the chances of exact match become too slim.
To address the challenges above, we make the following contributions:
@@
    We propose an end-to-end factual recall evaluation framework which is tailored to RAG systems. It can be used to (i) create a test dataset and (ii) perform automated factual recall evaluation given a RAG system (or more loosely an LM).
@@
    We probe into the performance of fact verification formulations in conditions of highly incomplete or inaccurate generated text. To do that, we augment WikiEval [^1] https://huggingface.co/datasets/
@@
  3. 3.
    We introduce Facts as a Function (FaaF), a new fact verification formulation which outperforms fact verification via prompting in all examined conditions and reduces the required number of LM calls and completion tokens by more than 5 times.
@@
    We open source the factual recall evaluation framework with FaaF as a python package (pip install faaf) and the augmented WikiEval dataset to help the community include factual recall in the RAG optimisation and ultimately build more reliable systems.
@@ 2 Related Work
Min et al. (2023) use prompt variations to directly verify individual facts extracted from LM generated biographies. They compare their method with human evaluation and find low error rates when retrieving the ground truth for the evaluated fact.
@@
Azaria and Mitchell (2023) also find that fact verification by prompting insufficient and propose to train a classifier on the hidden layer activations of open source LMs to predict the truthfulness of a generated statements. However, current leading commercial models (GPT-4) are lacking layer activation access and so require alternative methods.
@@
Fu et al. (2023) build on the concept of utilising token probabilities, introducing a self-evaluation framework for LMs. This framework leverages few-shot prompting to evaluate various instructed aspects of LM responses, such as factuallity, fluency, interest, among others.
@@
Zhang et al. (2020) attempt a flexible self-evaluation of generated text using reference answers (BertScore). BertScore calculates a similarity score between tokens in the generated and reference sentences using contextual embeddings. The key benefit being that there is no reliance on exact matching between generated and reference text. Nevertheless, a high semantic score at sentences level does not guarantee factual precision, especially when the information examined is not contextual and explicit (E.g. date).
@@
Kadavath et al. (2022) observing that LLMs offer well-calibrated probabilities for multiple-choice questions, effectively transform the validation of model-generated answers into a True/False multiple-choice question. This work offers an important insight in that constraining the the LM response into a binary option, it incites a sufficiently confident response which allows for self-evaluation.
Lastly, recent work on the the factual accuracy of LMs and RAG systems (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) took the approach of using the NaturalQuestions-Open (NQ-open) [^2] https://ai.google.com/research/NaturalQuestions dataset (Kwiatkowski et al., 2019) and calculate accuracy by judging whether any of the ground truth answers (NaturalQuestions annotations) appear in the predicted output via exact matching. NQ-open is a large scale dataset which comprises historical Google search queries and their human-annotated answers sourced from Wikipedia. Even though NQ-open, is valuable for its extensive scope and domain-agnostic nature, fact-verification via exact matches faces serious challenges. As highlighted by Cuconasu et al. (2024), a key issue is accurately determining the truthfulness of answers, especially when dealing with phrases that have the same meaning but (slightly different format or wording) or different date formats. Thus, the necessity for a more advanced analysis of answer variations is recognise and left for future research.
@@ 3 Facts as a Function
Facts as a Function (FaaF) introduces a streamlined fact verification method using structured objects for multi-fact assessment. Although in this paper we focus in examining the FaaF approach in the scope of factual recall evaluation in RAG systems, the method itself can equivalently be used for broader fact verification tasks.
Key idea 1: JSON over prompt
We propose that by using the function calling ability of the LM, we enforce a more formal mode of reasoning and communication compared to natural language for a number of reasons. First, by training on code and function definitions the LM learns the independence of the function variables with each other, this results in a clearer individual assessment of each requested variable. In addition, by leveraging function variable metadata and type hints, the LM can be better constrained to the accepted modes of response for each variable. This provides a more consistent guidance compared to instructions given once at the end or the beginning of a prompt. Building on the findings of Kadavath et al. (2022), who established that LMs show well calibrated probabilities when presented with multiple choice questions, we propose that using the function’s type definition of the variables to convey the accepted LM responses is a step further in the same direction, which aids the calibration of the token probabilities. Lastly, due to the strict nature of code syntax compared to natural language, gradients during training are expected to be steeper and thus the LM learns to do a better job at adhering to the expected output and respond with lower stochasticity - which is beneficial in the fact verification use case.
@@
As discussed above, a function definition can encapsulate a sufficient number of variables to be used by the LM so that they capture all the facts statements which need to be verified. Therefore, we move away from the concept that each fact should be verified individually via a fact-specific prompt and we propose instantiating a function per LM-generated text which needs to be factually assessed. The function can be programatically generated in a manner such that it includes all the individual facts which need to be verified in a given piece of text. In other terms, function calling allows for the fact verification of a long-form text as a unit, with a single LMeval call. This approach results in a reduction of cost and time for fact-verification which is multiple to the number of facts which would otherwise need to be assessed individually, as seen in Chen et al. (2022); Gao et al. (2023); Min et al. (2023); Lee et al. (2023).
@@
Using function definitions to communicate with the LM, enables access to a multitude of tools and further logic processing we can execute on LMeval’s output. This strategy permits us to delegate certain deterministic judgments away from LMeval. We demonstrate this capability by mapping a range of LM responses into a binary format (True / False). In doing so, the calibration of the LM response is enhanced as we provide a more accurate representation of the spectrum of potential outcomes than a simple True/False dichotomy. The underlying intuition is that ultimately, we can afford to ask LMeval simpler and clearer questions which can be answered more reliably and further process the LM output into a final response.
@@
This work introduces the idea of presenting the facts to the LM as a callable function. Let $S$ be the list of fact statements as strings to verify.
$S=[s_{1},s_{2},\ldots,s_{n}].$
A constructor function $C$ then maps the input list of facts $S$ and control parameters $P$ to an object $O$ with fields $f$ and an object-specific parsing function $G_{O}$.
$C(S,P)\to\left(O\left(f_{1},f_{2},\ldots,f_{n}\right),G_{O}\right).$
Each field from $(f_{1},f_{2},\ldots,f_{n})$ corresponds to a fact statement and is further parameterised by $P$. Control parameters $P$ include the methods and field properties which are injected into resulting object $O$. The object-specific parsing function $G_{O}$ is tasked with parsing the LM response in accordance to the schema and embedded instructions and methods found in $O$. Before passing object $O$ to a language model, we convert it to a JSON representation:
@@
Let $M$ be the language model used for fact verification (LMeval). The input of $M$ is a concatenation of $JSON_{O}$, a prompt $q$ which instructs $M$ to utilise $O$ and the input text $x$ which is to be assessed for factual recall with respect to the given facts $S$.
@@
$M$ responds with the output $o_{x}$ which is of string type. Then, the parsing function $G_{O}$ is used to convert the raw response $o_{x}$ to a callable function object $O^{\prime}$ which is an instance with values assigned on its fields.
$G_{O}(o_{x})\to\acute{O}$
@@
$\acute{O}=G_{O}(M(\text{JSON}_{O},q,x)).$
@@ 4 Assessment of fact verification formulations in the RAG setting
Figure 2 outlines the the factual recall evaluation framework which also serves as the experimental setup which we use to compare fact verification formulations with each other. Starting from the ground truth answer containing the desired information to fully answer the posed question, we derive a set of fact statements using a fact-generator LM (LMf). We then use these derived facts to evaluate each of the other answer variants of WikiEval for their factual recall via LMeval. In this manner, each answer is evaluated with respect to the information that is expected from it. It is easy to see how this framework could be applied in the RAG setting where different configurations or model choices impact the quality of the final response to the same input question and knowledge base.
@@
In order to probe into performance of automatic fact verification methods we chose to work with the WikiEval dataset (Es et al., 2023) which features three versions of answers, of variable quality, to questions on Wikipedia articles. Specifically, for each question, there is an answer (referred to as ground truth answer in this text for clarity), ungrounded answer and poor answer. All answers have been generated with GPT-3.5-turbo-16k.
The ground truth answer is generated by providing the LM with the correct context from the respective Wikipedia page. The ungrounded answer is generated by not providing any context to the LM. Finally, the poor answer is generated by instructing the LM to give an incomplete answer to the passed question. An example of each answer type can be seen in Figure 2. Generally, ungrounded answer answers feature false, incomplete and redundant information with respect to the ground truth answers. On the other hand, poor answers contain mostly incomplete information compared to ground truth answers i.e. no evidence for support or rejection for ground truth facts.
This dataset enables us to focus on the performance of the fact-verification as a function of the quality and completeness of different answer variants. In that way, we test closely the ability of different fact verification methods to identify unsupported facts when presented with (i) incorrect, (ii) indirectly relevant and (ii) incomplete information with some degree of distinction.
@@
To prepare the WikiEval dataset, we initially generate factual statements that fully capture the information from the ground truth answers, followed by manual annotation of a version of these facts for each answer variant. We use gpt-4-0125 to generate facts from the questions and ground truth answers from the WikiEval dataset (QA pair) using the following prompt:
@@
Fact generation via the prompt above results in a variable number of facts for each ground truth QA pair which depends on the length and information density in the processed ground truth answer. This process yielded 281 individual facts for each type of answer (thus 843 facts in total considering answer, ungrounded answer and poor answer) with an average of 5.6 fact statements generated for every QA pair. The prompt has been designed to ensure that the generated facts are complete sentences, understandable independently of each other or any external references.
Answer Type Human Accuracy
 Answer 100
 Ungrounded Answer 30.24
 Poor Answer 8.53
@@
We outsource the fact verification of the generated facts against the triplet of answers in WikiEval (answer, ungrounded answer and poor answer) to human evaluators. In this manner we build a ground truth evaluation for each answer type, enabling us to assess the effectiveness of automated fact-verification methods against it. The accuracy from the human fact verification can be seen in Table 1 where the deterioration of the ungrounded answer and poor answer relative to the ground truth answer is evident.
@@
For each set of fact statements which encapsulate a ground truth answer, we construct a facts-specific object which also contains a parsing method. Since the created object contains all the input facts as fields, we essentially process the set of facts as a unit. A sample of the JSON representation of a function object containing the first fact can be seen bellow:
@@
    ‘fact_0:{’description’:"Respond
    to the claim in the ’title’ by
    using one of the accepted Enum
    types",
@@
   Ψ’title’: ’It is clear in
    the given passage that Pope
    Benedict XVI became the head
    of the Catholic Church and
    sovereign of the Vatican
    City State on April 19, 2005’
   Ψ’type’: ‘string’},
@@
Each field includes metadata which can be used to pass instructions, type annotations and the fact statement to be verified itself. In addition to the function object, we pass the following prompt:
@@
The answer in the prompt is the input text we want to evaluate against the respective facts which have been previously derived by the ground truth answer in our dataset. After LMeval generates a response, we process it using the tailored parsing method previously created and instantiate the function object that was passed into LMeval, populating it with the generated values in their designated fields.
@@
A function object with fields which only accept True or False as a response from LMeval.
@@
A function object with fields which only accept True, False or Not clear from the given passage as a response from LMeval. In this scenario, the parser will map Not clear from the given passage to False when processing the LMeval output. This is an example of applying logic post LM generation. The intuition behind this configuration is that the rejection of a claim based on contradicting evidence is conceptually different to the rejection of a claim based on absence of evidence and we help the LMeval’s calibration by providing a clear option for each.
@@
In this instance we construct a function object with two fields for each input fact. One field for the factual evaluation and one field where we instruct the LMeval to generate an exact excerpt from the input text which directly supports the fact in question (citation field). We place the citation field prior to the factual evaluation field so that LMeval is made to first try and find a supporting citation from the input text before verifying the fact that is being assessed. Similarly, the intuition here is that by asking LMeval to search and retrieve a specific citation from the input text which supports a specific fact, it will result in a better calibrated verification of the respective fact statement.
@@
In this configuration we combine the two approaches outlined above to explore their combined effect. In detail, we construct the function object to include citation fields and factual evaluation fields for each fact and we define True, False or Not clear from the given passage as accepted responses.
@@
We use the latest commercially available model which support function calling gpt-4-0125. A comprehensive assessment of open source models which support function calling is left for future study as there are none available at this time but we will extend this study to more models as they become available. Also, the focus of this work is on the different formulation of facts during verification and not the assessment of specific LMs.
@@
We use Error Rate (ER) between the human fact-verification and the fact verification formulation as the main indicator of their reliability.

### 08 Apr 2024 v2

# FaaF: Facts as a Function for the evaluation of generated text

@@ Abstract
The demand for accurate and efficient verification of information in texts generated by large language models (LMs) is at an all-time high, but remains unresolved. Recent efforts have focused on extracting and verifying atomic facts from these texts via prompting LM evaluators. However, we demonstrate that this method of prompting is unreliable when faced with incomplete or inaccurate reference information. We introduce Facts as a Function (FaaF), a new approach to the fact verification task that leverages the function-calling capabilities of LMs. FaaF significantly enhances the ability of LMs to identify unsupported facts in texts, while also improving efficiency and significantly lowering costs compared to prompt-based methods. Additionally, we propose a framework for evaluating factual recall in Retrieval Augmented Generation (RAG) systems, which we employ to compare prompt-based and FaaF methods using various LMs under challenging conditions.
@@ 1 Introduction
The adoption and transformative impact of large language models (LMs) across industries are significantly driven by their application in knowledge-intensive tasks.
@@
Figure 1: An overview of FaaF, a constructor dynamically creates a function object from a set of fact statements. Function calling allows LMeval to verify all facts within a single call when provided with an input reference text. FaaF significantly reduces the error rate in identifying unsupported facts compared to prompting whilst reducing the number of LMeval calls and output tokens by more than 5 times.
@@
In these applications, Retrieval Augmented Generation (RAG) is often used to integrate out-of-training knowledge to the LM (Lee et al., 2019). This is achieved by dynamically fetching and incorporating relevant data from a trusted reference source prior LM text generation, resulting in more informed, accurate, and contextually rich outputs.
Considering the importance of factual accuracy of generated text in this setting, a significant body of work has focused on automated ways for obtaining it. At a high level, two distinct schools of thought emerge - shaped by the intended use case, the available test datasets and required scale. First, authors who address factual precision —the truthfulness of each statement in a generated text— in both RAG and non-augmented LM generation scenarios (Chen et al., 2022; Zhang et al., 2023; Gao et al., 2023; Lee et al., 2023; Min et al., 2023; Azaria and Mitchell, 2023; Yuan et al., 2021; Fu et al., 2023; Wang et al., 2023; Kadavath et al., 2022; Manakul et al., 2023). Given that a generated text may include both accurate and inaccurate statements, a common approach is to initially extract a number of fact statements from the generated text and verify them individually by prompting an LM evaluator model with appropriate references. Secondly, other studies (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) attempt to evaluate LMs and RAG systems in terms of exact matching any one of a set of predefined correct answers in the generated text.
There is an apparent lack of work on the practical measurement of factual recall—the extent to which all the information required to sufficiently answer a question is included in the generated text. Factual recall is a key performance metric, particularly in the RAG scenario, because it directly captures the performance of retrieval and generation simultaneously and closely reflects the central purpose of the system. Factual recall is at least as important as factual precision for RAG systems, where a response can be factually precise but given the wrong context may be irrelevant to the posed question.
Although the exact match method can be viewed as a fast and efficient form of binary recall score (where the existence of a single accepted answer in the generated text signals success), it faces serious challenges and limitations in the verification task.
@@
As recognised by Cuconasu et al. (2024), exact matching of ground truth answers in the generated text is prone to false negatives since the ground truth information might be present but phrased differently. Additionally, this approach cannot be applied when the grounding information is longer than a few words since the chances of exact match become very slim. Lastly, being binary renders it a relatively poor metric considering that there may be plural grounding information which need to be included in the generated text for a complete answer.
Further, current approaches relying on verification of each fact independently can be prohibitively costly in time and resources. Specifically, RAG systems include many moving parts (knowledge base, retrieval, prompt formulation, LM) and require substantial tuning (Es et al., 2023) therefore the efficiency and speed of the evaluation task is a requirement for practical usage.
To address the gap above, we make the following contributions:
@@
    We introduce Facts as a Function (FaaF), a new fact verification formulation which outperforms fact verification via prompting and reduces the required number of LM calls and completion tokens by more than 5 times.
@@
    We propose an end-to-end factual recall evaluation framework which is tailored to RAG systems. It can be used to (i) create a test dataset and (ii) perform automated factual recall evaluation given a RAG system (or generally a language model).
  3. 3.
    We probe into the performance of fact verification formulations in conditions of highly incomplete or inaccurate generated text. To achieve that, we augment WikiEval [^1] https://huggingface.co/datasets/
@@
    We open source FaaF [^2] https://github.com/vasiliskatr/faaf, the factual recall evaluation framework and the augmented WikiEval dataset (WikiEvalFacts) [^3] https://huggingface.co/datasets/Vaskatr/WikiEvalFacts to help the community include factual recall in the RAG optimisation and ultimately build more reliable systems.
@@ 2 Related Work
Min et al. (2023) use prompt variations and an aggregate non-parametric probability of the tokens in a fact statement to directly verify individual facts extracted from LM generated biographies. They compare their method with human evaluation and find low error rates when retrieving the ground truth for the evaluated fact.
@@
Azaria and Mitchell (2023) also find that fact verification by prompting insufficient and propose to train a classifier on the hidden layer activations of open source LMs to predict the truthfulness of a generated statements. However, current leading commercial models are lacking layer activation access and so require alternative methods.
@@
Fu et al. (2023) build on the concept of utilising token probabilities, introducing a self-evaluation framework for LMs. This framework leverages few-shot prompting to evaluate various instructed aspects of LM responses, such as factuality, fluency, interest, among others.
@@
Zhang et al. (2020) attempt a flexible self-evaluation of generated text using reference answers (BertScore). BertScore calculates a similarity score between tokens in the generated and reference sentences using contextual embeddings. The key benefit being that there is no reliance on exact matching between generated and reference text. Nevertheless, a high semantic score at sentences level does not guarantee factual precision, especially when the information examined is not contextual and depends only on a small number of tokens (E.g. date).
@@
Kadavath et al. (2022) observe that LLMs offer well-calibrated probabilities for self-evaluation via constraining the LM response into multiple-choice and True/False questions. This work highlights that simply requiring discrete response options prior to text generation can aid the response calibration by effectively narrowing the available distribution of next tokens — which would alternatively include many semantically overlapping paraphrases.
Lastly, recent work on the factual accuracy of LMs and RAG systems (Cuconasu et al., 2024; Liu et al., 2023; Kandpal et al., 2023; Mallen et al., 2023) took the approach of using the NaturalQuestions-Open (NQ-open) [^4] https://ai.google.com/research/NaturalQuestions dataset (Kwiatkowski et al., 2019) and calculate accuracy by judging whether any of the ground truth answers (NaturalQuestions annotations) appear in the generated text via exact matching. NQ-open is a large scale dataset which comprises historical Google search queries and their human-annotated answers sourced from Wikipedia. Even though NQ-open, is valuable for its extensive scope and domain-agnostic nature, fact-verification via exact matching faces serious challenges. As highlighted by Cuconasu et al. (2024), a key issue is accurately determining the truthfulness of answers, especially when dealing with phrases that have the same meaning but (slightly different format or wording) or different date formats. Thus, the necessity for a more advanced analysis of answer variations is recognised and left for future research.
@@ 3 Facts as a Function
Facts as a Function (FaaF) is a streamlined fact verification method using function calling for multi-fact assessment. Although in this paper we focus in examining FaaF in the scope of factual recall evaluation, the method itself can equivalently be used for factual precision and broader fact verification tasks.
Key idea 1: JSON and XML over prompt
We propose that by using the function calling ability of the LM, we enforce a more formal mode of token generation compared to natural language for three reasons. First, by leveraging the metadata of function arguments, type annotations and tailored instructions can effectively constrain the LM to the accepted modes of response. This provides an element of structured repetition of the instructions towards the LM (via the metadata attached to each function argument) and ultimately results in a more consistent guidance compared to instructions given once at the end or the beginning of a prompt. Secondly, with formalising the type annotation, we can avoid relying on exact matching to interpret the LM’s text response—which can prove detrimental as we demonstrate in the results of this paper. Type annotations can be combined with custom types to essentially convert a function argument into a classification result to a multiple-choice question. Building on the findings of Kadavath et al. (2022), who established that LMs show well calibrated probabilities when presented with multiple choice questions, we propose that using the function argument’s type annotations to convey the accepted LM responses is a step further in the same direction. Thirdly, due to the strict nature of code syntax compared to natural language, gradients during training are expected to be steeper and thus the LM learns to do a better job at adhering to the expected output and responding with lower stochasticity—which is particularly helpful in the fact verification use case.
@@
As discussed above, a function definition can encapsulate a sufficient number of arguments to be used by the LM so that they capture all the facts statements which need to be verified. Therefore, we move away from the concept that each fact should be verified individually via a fact-specific prompt and we propose instantiating a function per LM-generated text which needs to be factually assessed. The function can be programatically generated in a manner such that it includes all the individual facts which need to be verified in a given piece of text. In other terms, function calling allows for the fact verification of a long-form text as a unit, with a single LM call. This approach results in a reduction of cost and time for fact-verification which is proportional to the number of facts which would otherwise needed to be assessed individually, as seen in Chen et al. (2022); Gao et al. (2023); Min et al. (2023); Lee et al. (2023).
@@
Using function objects to communicate with the LM, enables access to a multitude of tools and further processing we can execute on LM’s output. This strategy permits us to delegate certain deterministic judgments away from the language model. We demonstrate this capability by mapping a range of LM responses into a binary format (True / False). In doing so, the calibration of the LM response is enhanced as we provide a more accurate representation of the spectrum of potential outcomes than a simple True/False dichotomy. The underlying intuition is that ultimately, we can afford to ask simpler and clearer questions to the LM which can be answered more reliably and further process the LM output into a final response.
@@
We aim to present the facts to the LM as a callable function. Let $S$ be the list of fact statements as strings to verify.
$S=[s_{1},s_{2},\ldots,s_{n}]$
A constructor function $C$ then maps the input list of facts $S$ and control parameters $P$ to an function object $O$ with arguments $f$.
$C(S,P)\to O\left(f_{1},f_{2},\ldots,f_{n}\right)$
Each argument from $(f_{1},f_{2},\ldots,f_{n})$ corresponds to a fact statement and is further parameterised by $P$. Control parameters $P$ include the methods, argument properties and metadata which are injected into resulting object $O$. Such methods can describe for example a desired post-processing step on the values in the arguments $(f_{1},f_{2},\ldots,f_{n})$. Before passing object $O$ to a language model, we convert it to a JSON or XML representation—depending on the LM’s function calling requirements:
@@
Let $M$ be the language model used for fact verification (LMeval). The input of $M$ is a concatenation of $JSON_{O}$, a prompt $q$ which instructs $M$ to utilise $O$ and the input text $x$ which is to be assessed for factuality with respect to the given facts $S$.
@@
$M$ responds with the output $o_{x}$ which is of string type. Then, a parsing function $G_{M}$ which adheres to particular response schema of $M$ is used to parse the raw response $o_{x}$ and invoke $O$ by assigning values on its arguments, yielding $\acute{O}$. The values being assigned to each of $(f_{1},f_{2},\ldots,f_{n})$ being the verification result of the underlying fact statement. Upon invocation of $O$, the function arguments undergo type validation to ensure that its schema and type annotations are respected.
$G_{M}(o_{x})\to\acute{O}$
@@
$\acute{O}=G_{M}(M(\text{JSON}_{O},q,x)).$
@@ 4 Assessment of fact verification formulations in the RAG setting
Figure 2 outlines the factual recall evaluation framework which also serves as the experimental setup which we use to compare fact verification formulations with each other. Starting from the ground truth answer containing the desired information to fully address the posed question, we derive a set of fact statements using a fact-generator LM (LMf). We then use these derived facts to evaluate each of the other answer variants in WikiEval for their factual recall via LMeval. In this manner, each answer is evaluated with respect to the information that is expected from it. It is easy to see how this framework could be applied in the RAG setting where different configurations or model choices impact the quality of the final response.
@@
In order to probe into performance of automatic fact verification methods we chose to work with the WikiEval dataset (Es et al., 2023) which features three versions of answers, of variable quality, to questions on Wikipedia articles. Specifically, for each question, there is an answer (referred to as ground truth answer in this paper for clarity), ungrounded answer and poor answer. All answers have been generated with GPT-3.5-turbo-16k.
The ground truth answer is generated by providing the LM with the correct context from the respective Wikipedia page. The ungrounded answer is generated by not providing any context to the LM. Finally, the poor answer is generated by instructing the LM to give an incomplete answer to the given question. An example of each answer type can be seen in Figure 2. Generally, ungrounded answers contain false, incomplete and redundant information with respect to the ground truth answers. On the other hand, poor answers contain primarily incomplete information compared to ground truth answers i.e. no evidence for support or rejection for ground truth facts.
This dataset enables us to asses the impact of quality and completeness of different answer variants to the performance of fact-verification. In that way, we test closely the ability of different fact verification methods and LMs to identify unsupported facts when presented with (i) incorrect, (ii) indirectly relevant and (ii) incomplete information with some degree of distinction.
@@
To prepare the WikiEval dataset, we initially generate fact statements that fully capture the information from the ground truth answers, followed by manual annotation of the generated facts for each answer variant. We call the resulting augmented dataset WikiEvalFacts. We use gpt-4-turbo to generate facts from the questions and ground truth answers from the WikiEval dataset (QA pair) using the following prompt:
@@
Fact generation via the prompt above results in a variable number of facts for each ground truth QA pair which depends on the length and information density in the processed ground truth answer. This process yielded 281 individual facts, which were annotated for each answer type (thus 843 annotated facts in total considering ground truth answer, ungrounded answer and poor answer) with an average of 5.6 fact statements generated for every QA pair. The prompt has been designed to ensure that the generated facts are complete sentences, understandable independently of each other or any external references.
Factual Accuracy
Answer Type Human Evaluation
 Ground Truth Answer 100
 Ungrounded Answer 30.6
 Poor Answer 8.5
@@
We outsource the fact verification of the generated facts against the triplet of answers in WikiEval (ground truth answer, ungrounded answer and poor answer) to human evaluators. In this manner we build a ground truth evaluation for each answer type, enabling us to assess the effectiveness of automated fact-verification methods against it. The accuracy from the human fact verification can be seen in Table 1 where the factual accuracy of ground truth answers is 100% since all the generated facts are True by design. The deterioration of the ungrounded answer and poor answer relative to the ground truth answer is evident.
@@
For each set of fact statements which encapsulate a ground truth answer, we construct a facts-specific function object and a parsing function. Since the created function object contains all the input facts as arguments, we perform verification to the set of facts as a unit. An example of the JSON representation of a function object containing the first fact can be seen bellow:
@@
    ‘fact_0:{
    ’description’:"It
    is clear from the passage
    that Pope Benedict XVI became
    the head of the Catholic Church
    and sovereign of the Vatican
    City State on April 19, 2005.
    Respond by using one of the accepted
    Enum types.",
@@
   Ψ’type’: ‘string’
@@
Each function argument includes metadata which can be used to pass instructions, type annotations and the fact statement to be verified itself. In addition to the function object, we pass the following prompt:
@@
The answer in the prompt is the input text we want to evaluate against the respective facts which have been previously derived by the ground truth answer in our dataset. After LMeval generates a response, the parsing function is used to invoke the function object by supplying the arguments parsed from the LMeval’s response.
@@
A function object with arguments which only accept True or False as a response from LMeval. As seen from the JSON example above, these are specified as custom type annotations (enum).
@@
A function object with arguments which only accept True, False or Not clear from the given passage as a response from LMeval. In this scenario, further processing inside the function object will map Not clear from the given passage to False after invocation. This is an example of applying a simple processing step on the LM output, post-generation. The intuition behind this configuration is that the rejection of a claim based on contradicting evidence is conceptually different to the rejection of a claim based on absence of evidence and we help the LMeval’s calibration by providing a clear response option for each.
@@
In this instance we construct a function object with two arguments for each input fact. One argument for the factual evaluation and one argument where we instruct the LMeval to generate an exact excerpt from the input text which directly supports the fact in question (i.e. citation). We place the citation argument prior to the factual evaluation argument so that LMeval is made to first try and find a supporting citation from the input text before verifying the fact that is being assessed. Similarly, the intuition here is that by asking LMeval to search and retrieve a specific citation from the input text which supports a specific fact, it will result in a better calibrated verification of the respective fact statement.
@@
In this configuration we combine the two approaches outlined above to explore their combined effect. In detail, we construct the function object to include citation arguments and we define True, False or Not clear from the given passage as accepted responses.
@@
We use the latest commercially available models which support function calling gpt-4-turbo, gpt-3.5-turbo, claude-3-opus and claude-3-sonnet. We also examined the recently released (mistral-Large) [^5] https://mistral.ai/news/mistral-large/ LM but it was excluded from the results in this paper due to its high failure rate of over 80% in some cases in generating appropriately formatted responses, rendering its results non-contributory to the discussion. It is important to note that we did not allow models to retry in case of a failed parsing of their response or failed invocation of the function object due to formatting. FaaF introduces strict constraints on the expected LM response and by permitting only one attempt, we also assess the LM’s proficiency in formatting their response, as well as verifying factual accuracy.
In addition, we highlight the crucial role of system prompt to model performance. A change in the system prompt can impact significantly the fact verification accuracy of the LM. We kept the system prompts for the GPT models unchanged but modified Claude models’ system prompts to incorporate a 1-shot example of simple function calling. This adjustment is a result of following the official function-calling recommendations of the two model families at the time of writing this paper — GPT uses JSON, while Claude uses XML. As the pace of developments in this space is fast, we expect further changes and improvements in the optimal interface between LMs and tools/functions.
Lastly, a comprehensive assessment of open source models which support function calling is left for future study as there are none widely established available at this time.
@@
This paper’s metrics exclusively evaluate language models (LMs) based on their successfully formatted responses where applicable (FaaF). In doing so, we ensure that the comparison of the LMs’ fact verification ability is not influenced by their capacity to format responses correctly — which is discussed separately.
@@
We use Error Rate (ER) between the human fact-verification and the fact verification formulation as the main indicator of verification accuracy.